OpenStudy (danysanchy97):

4/15 of the 315 members of a book club are male. How many female members are there in the club? A. 131 B. 231 C. 84 D. 174

OpenStudy (danysanchy97):

@rebeccaxhawaii @sweetburger

OpenStudy (ineedhelplz):

The answer is C. 84. Basically just multiple the two numbers to find 4/15 of 315. 4/15 equals 4 x 1/15. so, 4/15 x 315 = 4 x 1/15 x 315 which equals 4 x 315/15 which equals 4 x 21 which equals 84.

OpenStudy (ineedhelplz):

:)

OpenStudy (anonymous):

actually, that is a number of males in club. Number of females is 315-84 = 231

OpenStudy (isaidavila):

If there are 4/15 males then have to be 11/15 females 11/15*315= 231

OpenStudy (danysanchy97):

@KendrickLamar2014

OpenStudy (danysanchy97):

@Wendy.Ivette11714